In a class there are 20 boys and 15 girls. The ratio of boys to girls is.

a. 4:3

b. 3:4

c. 4:5

d. none of these


1
Expert's answer
2021-08-16T15:10:03-0400
"\dfrac{Boys}{Girls}=\dfrac{20}{15}=\dfrac{4}{3}"

The ratio of boys to girls is

a. 4:3